# Constrained Contextual Online Decision Making: A Unified Framework

### v1

@@ 4 Motivating Examples
To better illustrate the scope and relevance of our general decision-making model introduced in Section 3.1, we present several representative real-world applications that naturally fit within our framework. These include constrained bandits, stream-based active learning, sequential hypothesis testing, and online model calibration. Each of these problems, though arising in different domains and driven by distinct practical goals, shares a common structure of making sequential decisions under uncertainty and constraints. We show that, by appropriately specifying the utility and constraint functionals, our framework can unify and generalize existing approaches to these problems, thereby providing a systematic foundation for constrained sequential learning. Detailed formulations of these examples are provided in the remainder of this section.
4.1 Contextual Bandits with Safety Constraints
We consider the contextual bandits with safety constraint setting. At each round $t$, given context $x_{t}$, the agent selects an action $a_{t}\in\mathcal{A}$. The agent receives a reward cost pair $(r_{t},c_{t})$, where $r_{t}=R^{*}(x_{t},a_{t})+\xi_{t}^{r}$ and $c_{t}=C^{*}(x_{t},a_{t})+\xi_{t}^{c}$. Here, $R^{*}$ and $C^{*}$ are true expected reward and cost functions, and $\xi_{t}^{r}$, $\xi_{t}^{c}$ are i.i.d. zero-mean noise terms.
@@ 4.2 Active Learning with Budget Constraint
Following Amani et al. (2019), we assume a query cost $c(x,a)$ for each context-prediction pair. We define an augmented action as $(a,i)\in\mathcal{A}\times\left\{0,1\right\}$, where $i=1$ indicates querying the label. The corresponding augmented stochastic policy is $(\pi,p)$, where $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$ specifies the prediction distribution and $p:\mathcal{X}\times\mathcal{A}\rightarrow[0,1]$ gives the probability of querying $y$ given covariate prediction pair $(x,a)$. A loss function $l(a,y)$ quantifies prediction error. The utility distribution is given by the posterior label distribution $h^{*}(\cdot|x,a,i)=h^{*}(\cdot|x,a)$, because whether or not querying for the true label does not influence posterior distribution. The utility functional is defined as the negative expected loss:
${\mathcal{T}}_{1}(h^{*}(\cdot|x,a))=-\int_{y\in\mathcal{Y}}l(a,y)h^{*}(y|x,a)dy.$
We impose a per-round expected budget constraint. That is, for each $t\in[T]$, the augmented policy $(\pi_{t},p_{t})$ must satisfy $\sum_{a\in\mathcal{A}}c(x_{t},a)\pi_{t}(a|x_{t})p_{t}(x_{t},a)\leq\tau.$ Here, the constraint distribution for action $(a,i)$ is the delta distribution $\delta(c(x,a)\cdot\mathbb{I}\left(i=1\right))$, where $\mathbb{I}(x)$ stands for the indicator function. The constraint functional is thus:
${\mathcal{T}}_{2}(\delta(c(x,a)\mathbb{I}\left\{i=1\right\}))=c(x,a)\mathbb{I}\left\{i=1\right\},$
and taking the expectation under the augmented policy recovers the constraint expression.
@@ 4.3 Online Hypothesis Testing
We consider the online hypothesis testing in healthcare  (Bartroff et al., 2012; Jennison and Turnbull, 1999). A total of $T$ patients arriving sequentially, each associated with symptom $x_{t}$ sampled i.i.d. from $Q_{x}$. The hypothesis classes are =$\mathcal{H}=\left\{H_{0},H_{1}\right\}$, where $H_{0}$ corresponds to being healthy and $H_{1}$ stands for illness, respectively. After receiving $x_{t}$ at every round, the DM selects a diagnostic action $a_{t}\in\left\{0,1\right\}$, where 00 means healthy and $1$ means ill. The true label $w_{t}=\mathbb{I}\left\{x_{t}\in H_{1}\right\}$ is observed only after the action through some other expensive medical methods such as surgery. For any covariate $x$ and action $a$, the associated random variable is a Bernoulli random variable $w(x,a)\sim\text{Ber}(\mathbb{P}(w=a|x))$. We denote the probability mass function of this Bernoulli random variable as $f^{*}_{x,a}$. The DM aims to maximize the statistical power while controlling the level of critical mistakes under some specific level $\alpha$. Formally, the objective is to solve
@@
Therefore, for any $x,a$, the utility functional ${\mathcal{T}}_{1}$ is ${\mathcal{T}}_{1}(f^{*}_{x,a})=\mathbb{E}\left[\mathbb{I}\left\{w=a=0\right\}|x\right]$. Similarly, the constraint functional is ${\mathcal{T}}_{2}(f^{*}_{x,a})=\mathbb{E}\left[\mathbb{I}\left\{w=1,a=0\right\}|x\right]$.
@@ 4.4 Sequential $L_{1}$ Model Calibration
where $p_{t}$ is a function of $x_{t}$ and the prior history ${(p_{1},x_{1},y_{1}),\cdots,(p_{t-1},x_{t-1},y_{t-1})}$.
For each pair $(x,p)$, the associated probability distribution is the posterior distribution of $\left|q-\mathbb{I}{(y_{t}\leq p_{t})}\right|$. Denoted its pdf as $f(\cdot|x_{t},p_{t})$. The utility functional ${\mathcal{T}}_{1}$ is defined as the expectation of this quantity, i.e., ${\mathcal{T}}_{1}(f(\cdot|x_{t},p_{t}))=\int yf(y|x_{t},p_{t})dy$. and there is no constraint functional in this example.
5 Estimation Oracle &\&& Generalized Eluder Dimension
@@
Intuitively, the assumption 5.2 ensures that small estimation errors in probability distributions measured by the divergence $\mathsf{D}$ transform into proportionally small errors in the evaluations of utility and constraint functionals. This continuity property enables the decision maker to reliably approximate optimal decisions based on estimated distributions.
@@ 5.1 Density Estimation Oracle
Assuming that, for the given statistical functionals ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$, we have identified a metric-like divergence $\mathsf{D}$ under which both functionals are Lipschitz continuous, our next goal is a density estimation oracle that provides statistical guarantees with respect to $\mathsf{D}$.
@@ 5.2 Generalized Eluder Dimension
Lemma 5.6.
Lemma 5.7.
With the same condition as in lemma 5.6, with probability $1$,
6 Contextual Decision-Making with Constraints
@@ 6.1 Algorithm Design
We now present the pseudo-code of Algorithm 1 and explain its components in detail.
Algorithm 1 General Constrained Decision-Making
@@
     Sample action $a_{t}\sim\pi_{t}:=\tilde{\pi}_{t,t}$ and observe sample point $y_{t}$ from distribution $f^{*}_{x_{t},a_{t}}$ and $g^{*}_{x_{t},a_{t}}$.
@@ 7.1 Statistical Guarantees of the Algorithm
For any $\delta\in(0,1)$, with probability at least $1-\delta$, for any $t$, $\tilde{\Pi}_{\delta}^{t}(x)$ is a feasible policy set.
@@
For any $\delta\in(0,1)$, with probability at least $1-\delta$, setting
@@ 7.2 Extension to Infinite Function Class
If model classes $\mathcal{F}$ and $\mathcal{G}$ satisfy 7.7. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$, setting
@@ 10 Discussion and Future Work
Looking forward, our work opens up several promising directions. First, a natural extension is to incorporate state transitions and move toward Markov Decision Process (MDP) settings. This would enable the design of unified algorithms for general statistical tasks in MDPs, beyond the traditional reward maximization paradigm. Second, the generalized eluder dimension we introduce serves as a new lens for understanding the learnability of density model classes, and deserves further theoretical investigation. In particular, exploring how generalized eluder dimensions behave under different divergence measures for the same model class remains an open and intriguing question. Lastly, we believe it is valuable to apply our algorithm in real-world systems and empirically evaluate its practical performance across different domains.
@@ Appendix A Useful Math Tools
This lemma is called Azuma-Hoeffding inequality. For example, see [Sason, 2011].
@@
This is called Freedman’s inequality. See proof in Agarwal et al. [2012]. Then we could have lemma A.3.
@@
The proof of lemma A.4 could be found in Foster et al. [2021]. Finally, we invoke the following result called Weyl’s inequality about the changes to eigenvalues of a Hermitian matrix that is perturbed, [Franklin, 2012].
@@
Appendix B Proofs in Section 5
Proof of lemma 5.6.
@@ Appendix B Proofs in Section 5
By lemma 5.6, this can only occur when $t<(\frac{4r_{T}^{1/2}}{\varepsilon}+1)\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)$. Moreover, we have $\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)\leq\text{dim}_{E}(\mathcal{G},\mathsf{D},\alpha_{T})=d$ for $\varepsilon\geq\alpha_{T}$. Therefore, when $\omega_{i_{t}}>\varepsilon\geq\alpha_{T}$, $t\leq(\frac{4r_{T}^{1/2}}{\varepsilon}+1)$ and thus $\varepsilon\leq\sqrt{\frac{4r_{T}d}{t-d}}$. Therefore, if $\omega_{i_{t}}>\alpha_{T}$ we can conclude that $\omega_{i_{t}}\leq\min\left\{C,\sqrt{\frac{4r_{T}d}{t-d}}\right\}$. Then we have
@@ Appendix C Proofs in Section 7
Now we can prove the following important estimation error lemma.
Lemma C.2.
Consider a measurable randomized contextual bandit algorithm that selects stochastic policy $\pi_{t}$ based on $H_{t-1}$. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $t>K$ and every stochastic policy $\pi\in\Pi$, the expected estimation error is upper bounded by
Proof of lemma C.2.
@@
We first apply lemma C.2 and use the AM-GM inequality to get:
@@
From lemma C.2, we also know that with high probability, for any policy $\pi$
@@
If $\pi^{*}(x)\in\tilde{\Pi}^{t}_{\delta}(x)$, then we could directly apply lemma C.2 to get the result. Now we focus on the case where $\pi^{*}(x)\notin\tilde{\Pi}^{t}_{\delta}(x)$. Let $\pi_{0}(x)=a_{0}$ be the policy that always applies safe action. By definition, ${\mathcal{T}}_{2}(g^{*}_{x,a_{0}})=c_{0}$. Consider a mixed policy, given any $x$, $\tilde{\pi}_{t}(\cdot|x)=\gamma_{t}^{x}\pi^{*}_{t}(\cdot|x)+(1-\gamma_{t}^{x})\pi_{0}(\cdot|x)$. Thus, we define $\gamma_{t}^{x}$ to be the smallest number such that under this mixed policy, $\max_{g\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\tilde{\pi}_{t}})=\tau$. Recall that $\widetilde{V}^{t}_{\delta}(x,\pi)=\max_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})$, for the optimal stochastic policy $\pi^{*}$, it is obvious that $\pi^{*}(\cdot|x)$ is also optimal for every context $x$, let $\widetilde{g}^{x,t}=\mathop{\mathrm{argmax}}_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\pi^{*}})$.
@@
First, by lemma C.2, we have
@@
For the first term, again we apply lemma C.2,
@@
Specifically, we apply lemma 5.7 to get
@@
uniformly over all $t\geq 2$ and sequence $f^{2},f^{3},\cdots\in\mathcal{F}$. Then similar to the proof of lemma C.2, we could show that

### v3

@@ 4 Motivating Examples
To better illustrate the scope and relevance of our general decision-making model introduced in Section 3, we present several representative real-world applications that naturally fit within our framework. These include constrained bandits, stream-based active learning, sequential hypothesis testing, and online model calibration. Each of these problems, though arising in different domains and driven by distinct practical goals, shares a common structure of making sequential decisions under uncertainty and constraints. We show that, by appropriately specifying the utility and constraint functionals, our framework can unify and generalize existing approaches to these problems, thereby providing a systematic foundation for constrained sequential learning. Detailed formulations of these examples are provided in the remainder of this section.
4.1 Constrained Contextual Bandits
We consider the contextual bandits with the safety constraint setting. At each round $t$, given context $x_{t}$, the agent selects an action $a_{t}\in\mathcal{A}$. The agent receives a reward-cost pair $(r_{t},c_{t})$, where $r_{t}=R^{*}(x_{t},a_{t})+\xi_{t}^{r}$ and $c_{t}=C^{*}(x_{t},a_{t})+\xi_{t}^{c}$. Here, $R^{*}$ and $C^{*}$ are true expected reward and cost functions, and $\xi_{t}^{r}$, $\xi_{t}^{c}$ are i.i.d. zero-mean noise terms.
@@ 4.2 Active Learning with Budget Constraint
Following Amani et al. (2019), we assume a query cost $c(x,a)$ for each context-prediction pair. We define an augmented action as $(a,i)\in\mathcal{A}\times\left\{0,1\right\}$, where $i=1$ indicates querying the label. The corresponding augmented stochastic policy is $(\pi,p)$, where $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$ specifies the prediction distribution and $p:\mathcal{X}\times\mathcal{A}\rightarrow[0,1]$ gives the probability of querying $y$ given covariate prediction pair $(x,a)$. A loss function $l(a,y)$ quantifies prediction error.
In this example, we want to minimize the expected prediction error while maintaining low budget consumption. Specifically, we impose a per-round expected budget constraint. That is, for each $t\in[T]$, the augmented policy $(\pi_{t},p_{t})$ must satisfy $\sum_{a\in\mathcal{A}}c(x_{t},a)\pi_{t}(a|x_{t})p_{t}(x_{t},a)\leq\tau.$
[…]
@@ 4.3 Online Hypothesis Testing
We consider the online hypothesis testing in healthcare  (Bartroff et al., 2012; Jennison and Turnbull, 1999). A total of $T$ patients arriving sequentially, each associated with symptom $x_{t}$ sampled i.i.d. from $Q_{x}$. The hypothesis classes are =$\mathcal{H}=\left\{H_{0},H_{1}\right\}$, where $H_{0}$ corresponds to being healthy and $H_{1}$ stands for illness, respectively. After receiving $x_{t}$ at every round, the DM selects a diagnostic action $a_{t}\in\left\{0,1\right\}$, where 00 means healthy and $1$ means ill. The true label $y_{t}=\mathbb{I}\left\{x_{t}\in H_{1}\right\}$ is observed only after the action through some other expensive medical methods such as surgery. The DM aims to maximize the statistical power while controlling the level of critical mistakes under some specific level $\alpha$. Formally, the objective is to solve
@@
[…]
@@ 4.4 Sequential $L_{1}$ Model Calibration
where $p_{t}$ is a function of $x_{t}$ and the prior history ${(p_{1},x_{1},y_{1}),\cdots,(p_{t-1},x_{t-1},y_{t-1})}$. The goal of the DM is to minimize the expected calibration error.
[…]
5 Estimation Oracle and Generalized Eluder Dimension
@@
Intuitively, 5.2 ensures that small estimation errors in probability distributions measured by the divergence $\mathsf{D}$ transform into proportionally small errors in the evaluations of utility and constraint functionals. This continuity property enables the decision maker to reliably approximate optimal decisions based on estimated distributions.
@@ 5.1 Density Estimation Oracle
Assuming that, for the given functionals ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$, we have identified a metric-like divergence $\mathsf{D}$ under which both functionals are Lipschitz continuous, our next goal is a density estimation oracle that provides statistical guarantees with respect to $\mathsf{D}$.
@@ 5.2 Generalized Eluder Dimension
6 A Unified Algorithm: GED-UCB
@@ 6.1 Algorithm Design
We now present the pseudo-code of Algorithm 1 and explain its components in detail. We abbreviate our algorithm as GED-UCB, since it incorporates Generalized Eluder Dimension and Upper Confidence Bound design principle.
Algorithm 1 GED-UCB
@@
     Set $\pi_{t}:=\tilde{\pi}_{t,t}$.
     Sample action $a_{t}\sim\pi_{t}$ and observe sample point $y_{t}$ from distribution $f^{*}_{x_{t},a_{t}}$ and $g^{*}_{x_{t},a_{t}}$.
@@ 7.1 Statistical Guarantees of the Algorithm
For any $\delta\in(0,1)$, with probability at least $1-\delta$, for any $t$, and $x$, $\widetilde{\Pi}_{\delta}^{t}(x)$ is a feasible policy set.
@@
For any $\delta\in(0,1)$, with probability at least $1-\delta$, setting $\alpha_{r}=\frac{1-r_{0}}{\tau-c_{0}}$ and
@@ 7.2 Extension to Infinite Function Class
If model classes $\mathcal{F}$ and $\mathcal{G}$ satisfy 7.7. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$, setting $\alpha_{r}=\frac{1-r_{0}}{\tau-c_{0}}$ and
@@ 10 Discussion and Future Work
Looking forward, our work opens up several promising directions. First, a natural extension is to incorporate more complicated system dynamics such as Constrained Markov Decision Process and Partially Constrained Observable Markov Decision Process settings. This would enable the design of unified algorithms for more general statistical tasks. Second, the generalized eluder dimension we introduce serves as a new lens for understanding the learnability of density model classes, and deserves further theoretical investigation. In particular, exploring how generalized eluder dimensions behave under different divergence measures for the same model class remains an open and intriguing question. Lastly, we believe it is valuable to apply our algorithm in real-world systems and empirically evaluate its practical performance across different domains.
@@ Appendix A Useful Math Tools
This lemma is called Azuma-Hoeffding inequality. For example, see (Sason, 2011).
@@
This is called Freedman’s inequality. See proof in Agarwal et al. (2012). Then we could have lemma A.3.
@@
The proof of lemma A.4 could be found in Foster et al. (2021). Finally, we invoke the following result called Weyl’s inequality about the changes to eigenvalues of a Hermitian matrix that is perturbed, (Franklin, 2012).
@@
Appendix B Properties about Generalized Eluder Dimension
In this section, we provide two important lemmas about the generalized eluder dimension introduced in Section 5 and their proofs. They play an important role in proving the theorem 7.6.
Lemma B.1.
Lemma B.2.
With the same condition as in lemma B.1, with probability $1$,
Proof of lemma B.1.
@@ Appendix B Properties about Generalized Eluder Dimension
By lemma B.1, this can only occur when $t<(\frac{4r_{T}^{1/2}}{\varepsilon}+1)\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)$. Moreover, we have $\text{dim}_{E}(\mathcal{G},\mathsf{D},\varepsilon)\leq\text{dim}_{E}(\mathcal{G},\mathsf{D},\alpha_{T})=d$ for $\varepsilon\geq\alpha_{T}$. Therefore, when $\omega_{i_{t}}>\varepsilon\geq\alpha_{T}$, $t\leq(\frac{4r_{T}^{1/2}}{\varepsilon}+1)$ and thus $\varepsilon\leq\sqrt{\frac{4r_{T}d}{t-d}}$. Therefore, if $\omega_{i_{t}}>\alpha_{T}$ we can conclude that $\omega_{i_{t}}\leq\min\left\{C,\sqrt{\frac{4r_{T}d}{t-d}}\right\}$. Then we have
@@ Appendix C Proofs in Section 7
Now we can prove the important estimation error lemma 7.2.
Proof of lemma 7.2.
@@
We first apply lemma 7.2 and use the AM-GM inequality to get:
@@
From lemma 7.2, we also know that with high probability, for any policy $\pi$
@@
If $\pi^{*}(x)\in\tilde{\Pi}^{t}_{\delta}(x)$, then we could directly apply lemma 7.2 to get the result. Now we focus on the case where $\pi^{*}(x)\notin\tilde{\Pi}^{t}_{\delta}(x)$. Let $\pi_{0}(x)=a_{0}$ be the policy that always applies safe action. By definition, ${\mathcal{T}}_{2}(g^{*}_{x,a_{0}})=c_{0}$. Consider a mixed policy, given any $x$, $\tilde{\pi}_{t}(\cdot|x)=\gamma_{t}^{x}\pi^{*}_{t}(\cdot|x)+(1-\gamma_{t}^{x})\pi_{0}(\cdot|x)$. Thus, we define $\gamma_{t}^{x}$ to be the smallest number such that under this mixed policy, $\max_{g\in\tilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\tilde{\pi}_{t}})=\tau$. Recall that $\widetilde{V}^{t}_{\delta}(x,\pi)=\max_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\pi})$, for the optimal stochastic policy $\pi^{*}$, it is obvious that $\pi^{*}(\cdot|x)$ is also optimal for every context $x$, let $\widetilde{g}^{x,t}=\mathop{\mathrm{argmax}}_{g\in\widetilde{C}_{\mathcal{G}}(t,\delta)}\tilde{{\mathcal{T}}_{2}}(g_{x,\pi^{*}})$.
@@
First, by lemma 7.2, we have
@@
For the first term, again we apply lemma 7.2,
@@
Specifically, we apply lemma B.2 to get
@@
uniformly over all $t\geq 2$ and sequence $f^{2},f^{3},\cdots\in\mathcal{F}$. Then similar to the proof of lemma 7.2, we could show that